# Best approximation of a three-variable function by sum of one-variable coordinate functions

### v1

@@ Abstract
1Center for Mathematics and its Applications, Khazar University, Baku, Azerbaijan
@@ 1.1 Problem statement in a general form and previous results
Let $f:X\to\mathbb{R}$ be a continuous function in the $n$ variables, where $X=X_{1}\times X_{2}\times\dots X_{n}$ is the Cartesian product of the compact Hausdorff spaces $n$. Consider the approximation of $f$ by means of $g_{i},i=1,\dots,n$, where $g_{i}$ is a continuous function on $X_{i}$ for $i=1,\dots,n$. In other words, the problem asks for
$E(f,X)=\inf_{(g_{1},g_{2},...,g_{n})}||f(x_{1},x_{2},\dots,x_{n})-\sum_{i=1}^{n}g_{i}(x_{i})||_{C(X)}=?$ (1.1)
@@
In general, this problem is related to Hilbert’s 13th problem. Using A.S. Kronrod’s [16] results, A.N. Kolmogorov [14] showed that any function $f\in C([0,1]^{n})$ is representable as a superposition of continuous functions of three variables. Developing this method, V.I. Arnold [1], [2] proved that each continuous function of three variables can be represented by superposition of continuous functions of two variables. This is a negative answer to Hilbert’s conjecture. Further, A.N. Kolmogorov [15] proved the following remarkable, stronger theorem in a simple, elementary and elegant way:
Kolmogorov’s Theorem. There exist increasing functions $\phi_{pq}\in C([0,1])$, $p=1,...,n$, $q=1,...,2n+1$, such that an arbitrary given function
@@
Thus, every continuous function on the $n$ variables can be represented by superposition of continuous functions of one variable and the simplest function of two variables, namely, the sum of variables.
In addition to the problems of finding a precise expression for a given function in terms of a combination of functions of fewer variables, it is natural to consider the best approximation of functions of several variables by combinations of functions of fewer variables. These problems arise in various areas of mathematics, for example, in solving integral equations (see [5]), in solving functional equations (see [4]), in saving information (see [6]), in scaling matrices (see [8], [9]), etc. In space $C([0,1]^{n})$ this problem was first studied by S. P. Diliberto and E. G. Straus [6] (the problem was proposed by RAND Corporation for the most efficient distribution of data in the small memory of the computers of the day). The same problem was investigated by Yu. P. Ofman [20] for bounded bivariant functions, and for more general sets by D. E. Marshall and A. G. O’Farrell in [17]. Diliberto and Straus gave a ’leveling algorithm’ to find the approximation error (1.1) for $n=2$ and propose a theorem that will ’analogously’ hold for $n\geq 3$. However, G. Aumann in his work [3] has shown that while the proof in [6] was complete for case $n=2$, the generalization of the algorithm suggested for $n\geq 3$ was incorrect. A direct counterexample refuting the ”leveling algorithm” in $n\geq 3$ was given by V.A. Medvedev [18]. In the paper [3] G. Aumann showed a necessary and sufficient condition on the class of approximating functions that the ”leveling algorithm” works.
Another generalization approach for $n\geq 3$ to the problem starts with the paper [10] by M. Golomb, where he gave an important duality formula for the problem (1.1). Moreover, the works of G. Aumann and [17] by D.E. Marshall and A.G. O’Farrell have shown that the proof of the duality given in [10] also had a gap. Future, based on the results of S.Ya. Khavinson [13] and K.G. Navada [19], V. Ismailov [12] proved that the formula given by M. Golomb is valid in a stronger form.
@@
Note that for the two-dimensional case, under the condition $f_{xy}\geq 0$ T.J. Rivlin and R.J. Sibher [21] established that
@@
L. Flatto [7] gave a different proof of this formula in his work, where a generalization to the approximation of a $n$ variable function by $n-1$ variable functions was suggested under similar conditions. In case $n\geq 3$ (even with additional conditions), there is no finite algorithm for calculate the error $E(f,X)$. In this paper, under the conditions $\Delta_{x}\Delta_{y}f\geq 0$, $\Delta_{x}\Delta_{z}f\geq 0$, $\Delta_{y}\Delta_{z}f\geq 0$, we will prove a formula to calculate the error (1.1) for the three-variable continuous function on $[0,1]^{3}$.
@@ 1.2 Basic definitions and formulas
Let’s give some definitions that will be used throughout this paper. Most definitions in this subsection are taken from the paper [12] (also can be found in the monograph [11]).
@@
$\sum_{\pi_{j}(x_{i})=\xi}\lambda_{i}=0,\forall j,0\leq j\leq n\text{ and }\forall\xi\in X_{j}$ (1.2)
@@
The definition of a projection cycle goes back to the M.Golomb [10]. The above-given definition is an equivalent form of Golomb’s definition taken from [12].
Remark 1.1.
For $n=3$, the left-hand side of the equation (1.2) becomes the sum over planes parallel to one of the coordinate planes. Thus, if a projection-cycle has a point on a plane parallel to one of the coordinate planes, then it has at least one other point with an opposite sign on the same plane.
@@
When we are given a weak projection cycle $p$ and a nonzero vector $\lambda$ satisfying the equation (1.3), by $(p,\lambda)$ we mean the projection cycle-vector based on the points of $p$ corresponding to nonzero components of $\lambda$.
@@
A projection cycle is called a minimal projection cycle if it has no proper subset that is a projection cycle. A projection cycle-vector $(p,\lambda)$ is called minimal if the set of points of $p$ corresponding to nonzero components of $\lambda$ is a minimal projection cycle.
@@
$\sum|\lambda_{i}|=1$
Definition 1.5.
A set of points $p=(x_{1},x_{2},...,x_{m}),x_{i}\in X$ where the set $X$ is the Cartesian product $X=X_{1}\times X_{2}\times\dots X_{n}$, is called a weak projection cycle if there exist some real numbers $\lambda_{1},\lambda_{2},..\lambda_{m}$ such that, at least one $\lambda_{i}$ is nonzero and,
$\sum_{\pi_{j}(x_{i})=\xi}\lambda_{i}=0,\forall j,0\leq j\leq n\text{ and }\forall\xi\in X_{j}$ (1.3)
The difference between a projection cycle and a weak projection cycle is that the second may contain points with zero weight. Thus, any projection-cycle is a weak projection cycle, but the converse is not true in general. For example, the set of points $\{(0,0),(0,1),(1,0),(1,1),(2,2)\}$ in $\mathbb{R}^{2}$ is a weak projection cycle but not a projection cycle.
The concept of a minimal projection cycle and a weak projection cycle was defined by name loop and weak loop in [19] by K.G. Navada.
The following theorem expresses Golomb’s formula (proposed in [10]), in revisited and improved form, from [12] (See: Theorem 2.8 in [12] or Theorem 3.37 in [11]).
@@
However, it is not a calculable formula; the formula (1.4) can be improved by comparing orthonormal sums over distinct projection-cycles under conditions like [21] or [7] used in their works to get a formula. In this paper, we will find a formula for the case $n=3$ and $X=[0,1]^{3}$.
Remark 1.3.
For simplicity, we use the term ”projection cycle-vector” instead of ”projection cycle, vector pair” used in [12] and [11].
@@ 2 Three variable case
Let $\Omega=[0,1]^{3}$ denote the unit cube and $f:\Omega\rightarrow\mathbb{R}$ be a given continuous function on $3$ variables. Consider the problem:
$E(f,\Omega)=\inf||f(x,y,z)-\phi(x)-\psi(y)-\omega(z)||_{\infty}=?$
@@
Let, $\Delta_{(x,y,z)}f$ be defined as
$\Delta_{(x,y,z)}f(u,v,w)=f(u+x,v+y,w+z)-f(u,v,w)$
For simplicity, let’s denote
$\Delta_{x}=\Delta_{(x,0,0)},\Delta_{y}=\Delta_{(0,y,0)},\Delta_{z}=\Delta_{(0,0,z)}$
@@
$\begin{cases}\Delta_{x}\Delta_{y}f\geq 0\\ \Delta_{x}\Delta_{z}f\geq 0\\ \Delta_{y}\Delta_{z}f\geq 0\end{cases}$ (2.1)
@@
$\begin{cases}f_{xy}\geq 0\\ f_{yz}\geq 0\\ f_{xz}\geq 0\end{cases}$ (2.2)
Let, for any finite sets $S$, the set $X_{S}$ denote the set of all projection cycle-vectors, which, each point belongs to the set $S$ and $X_{S,min}$ denote the set of all minimal projection cycle-vectors $(p,\lambda)\in X_{S}$.
We obtain the following result.
Theorem 2.1.
Let, $f\in C(\Omega)$ be a function satisfying the relations (2.1), then there exist a finite set $U=U_{f}$ such that,
$E(f,\Omega)=\max_{(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z},min}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}$ (2.3)
where, $U_{x},U_{y},U_{z}$ are the projections of $U$ to the coordinate lines $Ox,Oy,Oz$ respectively.
Remark 2.1.
For $n=2$, above-mentioned papers [6],[7], [11] gives similar teorems. But, while the set $U$ was the corners of the unit square, which is independent of $f$ in the case $n=2$, $U$ will be dependent on the function $f$ for $n=3$.
Below, we will prove the Theorem $2.1$ and moreover will explicitly construct the set $U$.
We start with applying Golomb’s Formula (1.4) and get:
@@
where, the supremum runs over all minimal projection cycle-vectors $(p,\lambda)$ on the set $\Omega$.
To proceed further, we will search the supremum (2.4) as
$E(f,\Omega)=\sup_{(p,\lambda)}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\sup_{A,B,C}\left({\max_{(p,\lambda)\in X_{A\times B\times C}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}}\right)$ (2.5)
where the last supremum runs over all triples of finite sets $A,B,C\subset[0,1]$ Obviously, for any fixed $U$ and any projection cycle-vector $(p,\lambda)$, there exist at least one triple of sets $A,B,C\subset[0,1]$ with $U\subset A\times B\times C$ such that $(p,\lambda)\in X_{A\times B\times C}$. On the other hand choosing the sets $A,B,C$ finite ensures that the above maximum over $X_{A\times B\times C}$ is attained. Hence, the relation (2.5) is well-defined.
Before proving the Theorem $2.1$ we have to prove some facts. Later, we will see that to prove the Theorem 2.1 for the class of functions satisfying the conditions (2.2) is equivalent to proving it for functions satisfying the following stricter inequalities
$\begin{cases}\Delta_{x}\Delta_{y}f>0\\ \Delta_{x}\Delta_{z}f>0\\ \Delta_{y}\Delta_{z}f>0\end{cases}$ (2.6)
Thus, below here we will work with functions satisfying (2.6).
@@ 2.1 Optimal Projection Cycle-Vectors on a Finite Set
First, we will investigate some properties of projection cycle-vectors that give the following maximum
$\max_{(p,\lambda)\in X_{A\times B\times C}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$ (2.7)
where $A,B,C\subset[0,1]$ are arbitrary chosen sets with $\{0,1\}\subset A\cap B\cap C$.
Let’s give some further definitions and auxiliary facts.
@@
$\xi_{1}\geq\xi_{2},\gamma_{1}\geq\gamma_{2},\zeta_{1}\geq\zeta_{2}\text{ or }\xi_{1}\leq\xi_{2},\gamma_{1}\leq\gamma_{2},\zeta_{1}\leq\zeta_{2}.$
@@
If two points$(\xi_{1},\gamma_{1},\zeta_{1})$ and $(\xi_{2},\gamma_{2},\zeta_{2})$ are not well-ordered, then, without loss of generality, we can assume that:
@@
Now we present an important lemma, which is a direct consequence of (2.6).
@@
holds, where we denote $(a,b)=\min(a,b)$ and $[a,b]=\max(a,b)$. The equality happens iff $x_{1}$ and $x_{2}$ are well-ordered.
@@
If $x_{1}$ and $x_{2}$ are well-ordered, then the expressions on the left and right-hand sides become the same, and the equality holds.
@@
$=\left(f(x_{2}^{\prime})+f(x_{2}^{\prime\prime})-f(x_{1}^{\prime\prime})-f(x_{2})\right)+\left(\left(f(x_{1}^{\prime\prime})+f(x_{1}^{\prime})-f(x_{2}^{\prime\prime})-f(x_{1})\right)\right)>0$
@@
$x$$z$$y$$x_{2}^{\prime\prime}$$x_{1}^{\prime\prime}$$x_{2}^{\prime}$$x_{2}$$x_{1}$$x_{1}^{\prime}$
@@
A projection cycle-vector $(p^{0},\lambda^{0})\in X_{S}$ is called optimal for the function $f$ on the finite set $S$ if
$\frac{\sum\lambda^{0}_{i}f(x^{0}_{i})}{\sum|\lambda^{0}_{i}|}=\max_{(p,\lambda)\in X_{A\times B\times C}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}$
The set of all optimal projection cycle-vectors for $f$ on $S$ is denoted by $X_{S,f}$.
Obviously, for any $(p,\lambda)\in X_{S,f}$ we have
$\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}\geq 0$ (2.9)
as otherwise, taking $(p,-\lambda)$ will contradict the optimality of $(p,\lambda)$.
@@
The sum of two projection cycles-vectors is a projection cycle-vector on the union of the points of the given projection cycle-vectors, with weight at each point equal to the sum of the weights of the given projection cycle-vectors on that point.
We can express the above definition by formulas as follows: $(p^{1},\lambda^{1})+(p^{2},\lambda^{2})=(p,\lambda)$ where, $p=p^{1}\cup p^{2}$ and
$\lambda(x)=\begin{cases}\lambda^{1}(x),\hskip 45.0ptx\in p^{1}\backslash p^{2}\\ \lambda^{1}(x)+\lambda^{2}(x),\hskip 8.0ptx\in p^{1}\cap p^{2}\\ \lambda^{2}(x),\hskip 43.0ptx\in p^{2}\backslash p^{1}\\ \end{cases}$ (2.10)
for any two (weak) projection cycle vectors $(p^{1},\lambda^{1})$ and $(p^{2},\lambda^{2})$.
Clearly,if both $(p^{1},\lambda^{1})\in X_{S}$ and $(p^{2},\lambda^{2})\in X_{S}$ for some finite set $S$ and function $f$, then
$(p^{1},\lambda^{1})+(p^{2},\lambda^{2})\in X_{S}$
since $p^{1}\subset S$ and $p^{2}\subset S$ is equivalent to $p^{1}\cup p^{2}\subset S$. Moreover, we have
$\sum_{x_{i}\in p}\lambda_{i}f(x_{i})=\sum_{x_{j}\in p^{1}}\lambda^{1}_{j}f(x_{j})+\sum_{x_{k}\in p^{2}}\lambda^{2}_{k}f(x_{k})$
@@
Let $(p^{1},\lambda^{1})$, and $(p^{2},\lambda_{\varepsilon})$ be two projection cycle vectors, where, $\lambda_{\varepsilon}$ is a vector of length $m$ such that each component is $\varepsilon$ or $-\varepsilon$ for some positive real number $\varepsilon$. Let $k$ be the number of common points of $p^{1}$ and $p^{2}$, such that the corresponding components of the vectors $\lambda$ and $\lambda_{\varepsilon}$ have opposite signs. If $k\geq\frac{m}{2}$, then for
@@
$\sum_{x\in p}|\lambda(x)|\leq\sum_{x\in p^{1}}|\lambda^{1}(x)|$ (2.11)
where, $(p,\lambda)=(p^{1},\lambda^{1})+(p^{2},\lambda_{\varepsilon})$.
@@
By the definition (2.10) we have,
$\sum_{x\in p}|\lambda(x)|-\sum_{x\in p^{1}}|\lambda^{1}(x)|=\sum_{x\in p^{1}\cap p^{2}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right)+\sum_{x\in p^{2}\backslash p^{1}}\varepsilon$
The first summand can be written as,
@@
$=\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)<0\end{subarray}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right)+\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)>0\end{subarray}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right)$

### v2

@@ Abstract
Keywords: Best approximation; coordinate functions; projection cycle; Golomb’s formula; minimal projection cycle.
2020 Mathematics Subject Classification: 41A50; 41A30
@@ 1.1 Problem statement in a general form and previous results
Let $f:X\to\mathbb{R}$ be a continuous function of $n$ variables, where $X=X_{1}\times X_{2}\times\dots X_{n}$ is the Cartesian product of $n$ compact Hausdorff spaces. Consider the approximation of $f$ by sums of $g_{i},i=1,\dots,n$, where $g_{i}$ is a continuous function on $X_{i}$ for $i=1,\dots,n$. In other words, the problem asks for
[…]
@@
In general, this problem is related to Hilbert’s 13th problem. Using A.S.Kronrod’s [27] results, A.N. Kolmogorov [25] showed that any function $f\in C([0,1]^{n})$ is representable as a superposition of continuous functions of three variables. Developing this method, V.I.Arnold [2], [3] proved that each continuous function of three variables can be represented by a superposition of continuous functions of two variables. This is a negative answer to Hilbert’s conjecture. Further, A.N.Kolmogorov [26] proved the following remarkable, stronger theorem in a simple, elementary and elegant way:
Kolmogorov’s Theorem. There exist increasing functions $\phi_{pq}\in C([0,1])$, p=1,,np=1,...,n, q=1,,2n+1q=1,...,2n+1, such that an arbitrary given function
@@
Thus, every continuous function on the $n$ variables can be represented by a superposition of continuous functions of one variable and the simplest function of two variables, namely, the sum of variables.
In addition to the problems of finding a precise expression for a given function in terms of a combination of functions of fewer variables, it is natural to consider the best approximation of functions of several variables by combinations of functions of fewer variables. These problems arise in various areas of mathematics, for example, in solving integral equations (see [10]), functional equations (see [9]), in saving information (see [11]), in scaling matrices (see [14], [15]), in isogeometric analysis (see [8]), etc.
In the space $C([0,1]^{n})$, this problem was first studied by S. P. Diliberto and E. G. Straus [11] (the problem was proposed by RAND Corporation for the most efficient distribution of data in the small memory of the computers of the day). The same problem was investigated by Yu. P. Ofman [33] for bounded bivariant functions, and for more general sets by D. E. Marshall and A. G. O’Farrell in [28]. In [11], Diliberto and Straus gave a “leveling algorithm” to find the approximation error (1.1) for $n=2$. The addressed problem, together with the methods developed from its analysis, finds important applications in approximation by subalgebras and, in particular, in RBF approximation, ridge function approximation, and neural network models. Recent papers (see [1, 4, 5, 6, 18]) improve the two-dimensional theory in connection with these directions. Moreover, based on the “levelling” algorithm, the ADI (alternating-direction-implicit) iteration method for numerical solution of differential equations is developed (see [38, 39, 40]).
In the same paper, Diliberto and Straus stated that a similar algorithm will work for $n\geq 3$. However, G. Aumann in his work [7] has shown that while the proof in [11] was complete for case $n=2$, the generalization of the algorithm suggested for $n\geq 3$ was incorrect and proved a necessary and sufficient condition on the class of approximating functions that the “leveling algorithm” works. A direct counterexample refuting the “leveling algorithm” in $n\geq 3$ was given by V.A. Medvedev [29]. This example shows that the levelling algorithm, nor its small variations, can converge to a best approximation from the sum of three subspaces (see [34]).
Another solution approach for $n\geq 3$ to the problem starts with the paper [16] by M. Golomb, where an important duality formula involving functionals defined on projection-cycles for the problem (1.1) was stated. However, later in [28] by D.E. Marshall and A.G. O’Farrell, it was shown that the proof of the formula given in [16] also has a gap. In [28], using the ergodic method, the lightning-bolt principle is also proven, which corresponds to the two-dimensional version of Golomb’s formula. It is worth remarking that the results obtained in the same paper in the two-dimensional case do not extend to sums such as $f(x)+g(y)+h(z)$ on subsets of $\mathbb{R}^{3}$. There is no way to generate the annihilators of a sum of three algebras using any analogue of the lightning bolts. This is related to the fact that ergodic theory is essentially limited to actions of groups that have subexponential growth. This shows that there are essentially new phenomena that arise with the sum of three subalgebras of $C(X)$, as opposed to two (see [32]).
Future, with the help of the results of S.Ya. Khavinson [24] and K.G. Navada [31], V. Ismailov [23] proved that the formula given by M. Golomb is valid in a stronger form. Although this formula itself involves finding a supremum over an infinite number of functionals and there is no closed formula to find the error explicitly, even with no additional conditions.
@@
The first finite formula for the two-dimensional case is established by T.J.Rivlin and R.J.Sibner in the paper [36], under the condition $f_{xy}\geq 0$,
@@
L. Flatto gave a different proof of the last formula in his work [12], where a general result for the approximation of an $n$ variable function by n1n-1 variable functions was suggested under similar conditions. Some other notable improvements in this direction have been made in [19] [20], where generalisations of the above formula in stair-like polygons have been proven. However, no finite formula has been proven for other domains, even for the simplest ones as the circle and triangle.
In case $n\geq 3$ (even with additional conditions), there is no finite formula and converging algorithm to calculate the error $E(f,X)$. In the present paper, under conditions $\Delta_{x}\Delta_{y}f\geq 0$, $\Delta_{x}\Delta_{z}f\geq 0$, $\Delta_{y}\Delta_{z}f\geq 0$, we will prove a finite formula to calculate error (1.1) in $0,1]^{3}$ (See Theorem 2.2).
The obtained results and methods proposed in this paper can be used in or extended to approximation by subalgebras and, in particular, in RBF approximation, ridge function approximation, and neural network models in the three-dimensional case. Moreover, the same ideas could lead to the construction of a three-dimensional levelling-type algorithm to find the best approximation functions, too. Such an algorithm would allow the construction of more efficient versions of existing numerical solution methods. As an example, the three-dimensional ADI iteration method itself proposed in [39] uses the two-dimensional levelling algorithm due to the lack of a converging three-dimensional alternative.
@@ 1.2 Basic definitions and formulas
We start with some definitions that will be used throughout this paper. Most of the definitions in this subsection are taken from the paper [23] (also can be found in the monograph [20]).
@@
$\sum_{\pi_{j}(x_{i})=\xi}\lambda_{i}=0,\ \forall\xi\in X_{j}\ \text{ and }\ \forall j,\ 0\leq j\leq n,$ (1.2)
@@
The definition of a projection cycle goes back to M.Golomb [16]. The above-given definition is an equivalent form of the Golomb’s definition taken from [23].
@@
Remark 1.1.
For $n=3$, the left-hand side of equation (1.2) becomes the sum over planes parallel to one of the coordinate planes. Thus, if a projection-cycle has a point on a plane parallel to one of the coordinate planes, then it has at least one other point with an opposite signed weight on the same plane.
@@
A projection cycle is called a minimal projection cycle if it has no proper subset that is a projection cycle. A projection cycle-vector $(p,\lambda)$ is called a minimal projection cycle-vector if the set of points of $p$ corresponding to nonzero components of $\lambda$ is a minimal projection cycle.
@@
$\sum|\lambda_{i}|=1.$
The concept of the minimal projection cycle was defined by name loop in [31] by K.G. Navada.
The following theorem expresses Golomb’s formula (proposed in [16]), in a revisited and improved form, from [23] (See: Theorem 2.8 in [23] or Theorem 3.37 in [20]).
@@
For simplicity, we use the term “projection cycle-vector” instead of “projection cycle, vector pair” used in [23] and [20].
In what follows, the following notations will be used. As usual $\Delta_{(x,y,z)}f$ is defined as
$\Delta_{(x,y,z)}f(u,v,w)=f(u+x,v+y,w+z)-f(u,v,w).$
For simplicity, we denote
$\Delta_{x}=\Delta_{(x,0,0)},\Delta_{y}=\Delta_{(0,y,0)},\Delta_{z}=\Delta_{(0,0,z)}.$
For any finite sets $S$, the set $[S]$ denotes the set of all projection cycle-vectors, which, each point belongs to the set $S$ and $S]_{min}$ denote the set of all minimal projection cycle-vectors $(p,\lambda)\in[S]$.
@@ 2 Three variable case
Let $\Omega=[0,1]^{3}$ denote the unit cube, and let $f:\Omega\rightarrow\mathbb{R}$ be a given continuous function of three variables. Consider the problem:
[…]
@@
$\begin{cases}\Delta_{x}\Delta_{y}f\geq 0\\ \Delta_{x}\Delta_{z}f\geq 0\\ \Delta_{y}\Delta_{z}f\geq 0.\end{cases}$ (2.1)
@@
$\begin{cases}f_{xy}\geq 0\\ f_{yz}\geq 0\\ f_{xz}\geq 0.\end{cases}$ (2.2)
We start by applying Golomb’s Formula (1.3) and get:
@@
where the supremum runs over all minimal projection cycle-vectors $(p,\lambda)$ on the set $\Omega$. To proceed further, we will search the supremum (2.3) as
$E(f,\Omega)=\sup_{(p,\lambda)}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\sup_{A,B,C}\left({\max_{(p,\lambda)\in[A\times B\times C]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}}\right),$ (2.4)
where the last supremum runs over all triples of finite sets $A,B,C\subset[0,1]$. Obviously, for any projection cycle-vector $(p,\lambda)$, there exists at least one triple of sets $A,B,C\subset[0,1]$ such that $(p,\lambda)\in[A\times B\times C]$. On the other hand, choosing the sets $A,B,C$ finite ensures that the above maximum over $A\times B\times C]$ is attained. Hence, the relation (2.4) is well-defined.
Before proving our main results, we have to prove some facts. From now on will proceed with the following stricter version of the inequalities (2.1)
$\begin{cases}\Delta_{x}\Delta_{y}f>0\\ \Delta_{x}\Delta_{z}f>0\\ \Delta_{y}\Delta_{z}f>0.\end{cases}$ (2.5)
Later, we will see that we can pass back to the original condition (2.1).
@@ 2.1 Optimal Projection Cycle-Vectors on a Finite Set
We start by investigating properties of projection cycle-vectors that give the following maximum
$\max_{(p,\lambda)\in[A\times B\times C]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|},$ (2.6)
where $A,B,C\subset[0,1]$ are the sets chosen arbitrary with $\{0,1\}\subset A\cap B\cap C$ and $f\in C(\Omega)$ is a function satisfying the conditions (2.1) or (2.5).
Let us give some further definitions and auxiliary facts.
@@
[…]
@@
[…]
@@
Now we present an important lemma, which is a direct consequence of the inequalities (2.5).
@@
holds, where we denote $(a,b)=\min(a,b)$ and $a,b]=\max(a,b)$. The equality happens iff $x_{1}$ and $x_{2}$ are well-ordered.
@@
If $x_{1}$ and $x_{2}$ are well-ordered, then the expressions on the left- and right-hand sides become the same, and the equality holds.
@@
$\left(f(x_{2}^{\prime})+f(x_{2}^{\prime\prime})-f(x_{1}^{\prime\prime})-f(x_{2})\right)+\left(\left(f(x_{1}^{\prime\prime})+f(x_{1}^{\prime})-f(x_{2}^{\prime\prime})-f(x_{1})\right)\right)>0,$
@@
[…]
@@
A projection cycle-vector $(p^{0},\lambda^{0})\in[S]$ is called optimal for the function $f$ on the finite set $S$ if
$\frac{\sum\lambda^{0}_{i}f(x^{0}_{i})}{\sum|\lambda^{0}_{i}|}=\max_{(p,\lambda)\in[A\times B\times C]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$
The set of all optimal projection cycle-vectors for $f$ on $S$ is denoted by $S]_{f}$.
Obviously, for any $(p,\lambda)\in[S]_{f}$ we have
$\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}\geq 0,$ (2.8)
as otherwise, taking $(p,-\lambda)$ will contradict the optimality of $(p,\lambda)$. And, moreover since $[S]$ is finite, $S]_{f}$ is always non-empty.
@@
The sum of two projection cycles-vectors is a projection cycle-vector on the union of the points of the given projection cycle-vectors, with weight at each point equal to the sum of the weights of the given projection cycle-vectors on that point. If we encounter any points with a weight of zero, we would ignore them.
We can express the above definition by formulas as follows:
$(p^{1},\lambda^{1})+(p^{2},\lambda^{2})=(p,\lambda),$
where $\pp=\left(p^{1}\cup p^{2}\right)\backslash p^{*}$, $p^{*}=\{q:q\in p^{1}\cap p^{2},\lambda^{1}(q)+\lambda^{2}(q)=0\}$ and
[…]
for any two projection cycle-vectors $(p^{1},\lambda^{1})$ and $(p^{2},\lambda^{2})$.
Clearly, if both $(p^{1},\lambda^{1})\in[S]$ and $(p^{2},\lambda^{2})\in[S]$ for some finite set $S$, then
$(p^{1},\lambda^{1})+(p^{2},\lambda^{2})\in[S]$
since $p^{1}\subset S$ and $p^{2}\subset S$ is equivalent to $p^{1}\cup p^{2}\subset S$. Moreover for any function $f$ defined on $0,1]^{3}$, we have
$\sum_{x_{i}\in p}\lambda_{i}f(x_{i})=\sum_{x_{j}\in p^{1}}\lambda^{1}_{j}f(x_{j})+\sum_{x_{k}\in p^{2}}\lambda^{2}_{k}f(x_{k}).$
In fact, we will never encounter a situation of adding two projection cycle-vectors of the form $(p,\lambda)$ and $(p,-\lambda)$. However, for the sake of correctness, we can still denote this sum as a vanishing projection cycle-vector. That is a projection cycle, with no points and no weights, and will not be considered an element of $[S]$, for any set $S$.
The following lemma demonstrates our way of constructing a “more optimal” projection-cycle vector from the given “non-optimal” one.
@@
Let $(p^{1},\lambda^{1})$ and $(p^{2},\lambda_{\varepsilon})$ be two projection cycle vectors, where $\lambda_{\varepsilon}$ is a vector of length $m$ such that each component is $\varepsilon$ or $\varepsilon$ for some positive real number $\varepsilon$. Let $k$ be the number of common points of $p^{1}$ and $p^{2}$, such that the corresponding components of the vectors $\lambda^{1}$ and $\lambda_{\varepsilon}$ have opposite signs. If $k\geq\frac{m}{2}$, then for
@@
$\sum_{x\in p}|\lambda(x)|\leq\sum_{x\in p^{1}}|\lambda^{1}(x)|,$ (2.10)
where $(p,\lambda)=(p^{1},\lambda^{1})+(p^{2},\lambda_{\varepsilon})$.
@@
By the equation (2.9) we have
$\sum_{x\in p}|\lambda(x)|-\sum_{x\in p^{1}}|\lambda^{1}(x)|=\sum_{x\in p^{1}\cap p^{2}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right)+\sum_{x\in p^{2}\backslash p^{1}}\varepsilon.$
The first summand can be written as
@@
$\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)<0\end{subarray}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right)+\sum_{\begin{subarray}{c}x\in p^{1}\cap p^{2}\\ \lambda_{1}(x)\lambda_{\varepsilon}(x)>0\end{subarray}}\left(|\lambda^{1}(x)+\lambda_{\varepsilon}(x)|-|\lambda^{1}(x)|\right).$